Laser power is reading zero

A typical laser operates with an output power between 300 and 500 mW, or up to 3 W (3000 mW) for a high-power laser. The laser power reading can be displayed on the status bar at the bottom of the PROspect window by enabling this feature in the Laser setup dialog (Calibration / Laser setup…) or viewed in the Advanced laser options dialog (Calibration / Laser setup… / Advanced). If the measured laser power is lower than expected, this can prevent the data acquisition process from continuing in some cases.

The following troubleshooting steps may help determine why the laser power is reading lower than expected.

  1. Check that the laser key is turned on and the power knob is turned to Max if applicable.
  2. Check for laser light coming from the optical probe by holding a white card in front of the end of the probe. A red dot should be visible for most lasers. Never look directly into the end of an optical fiber or probe! If light is visible but the software is still reporting 0 mW, see the section on Serial control board not responding.
  3. Check if the laser is enabled in the software. Navigate to Calibration / Laser setup… / Advanced… and ensure that there is a green indicator saying Enabled. If the laser is Disabled, enable it by pushing the Enable Laser A/B button.
  4. Make a note of the Set Power setting and the Power, Temp, and Current readings. Some older instruments will not display the laser Temp or Current. A typical laser will operate at 21 deg C. The laser-diode current will not turn on until the temperature has dropped below approximately 25 deg C. If the temperature never goes below 25 degrees, this would explain the laser power reading of zero. Most likely the laser should be returned for repair.
  5. If none of the laser readings appear correct, there might be a problem with the control board. See the section entitled Serial control board not responding.
  6. Continue to troubleshoot the laser itself by disconnecting it from the PI-200 instrument. Remove the lid of the instrument by either removing the thumbscrews for the rack-mounted PI-200I, or using a 5/64” hex key to remove the lid of the PI-200 lab instrument.
  7. Check that the laser itself has a key inserted and turned on, and that the green Laser On indicator is lit. Also check that the round power connector cable is attached. If the laser appears to be powered on correctly, but the Laser On indicator is not lit, the power supply in the laser may be out. Contact Process Instruments about returning the laser for repair.
  8. Check the power adjust knob on the laser itself. If it is not at Max, turn it up to Max and see if any light is now coming out of the probe. If so, you may want to try Adjusting the internal laser jumpers for your particular setup.
  9. To continue troubleshooting, check that the laser key is turned off on both the PI-200 instrument and on the laser itself. Unscrew the small fiber patch cable from the laser at the SMA connector on the laser itself. It is important that the fiber is not removed when the laser is powered on as this could damage the fiber tip! Turn on both laser keys and make sure the power knobs on the laser itself and on the instrument are at Max. If light is now visible then there is a problem with the fiber patch cable or the probe itself. Please Contact Process Instruments for further help.
  10. Now check if the laser works separate from the instrument. Again turn off both laser keys: on the instrument and on the laser itself. Remove the 15 pin control cable which connects the laser to the instrument control board. Open the top of the laser by removing the lid using a 5/64” hex key. Inside the laser case directly adjacent to the 15-pin plug, there are two jumpers which control the laser modulation and laser enable functions. Each can be set to either internal or external. In a typical PI-200 system, both of these functions are set to external. (The exception to this is a PI-200I rack mount system with the older style 9-pin serial control board. In this case the laser enable is set to external, but the laser modulation is set to internal.) Using a small needle-nose pliers, move both jumpers to their internal setting (INTMOD and INT LE). Now turn on both laser keys and ensure the power know on the laser is at Max. You should see light coming out of the fiber coupler after a 10-15 second warm-up period. If not, please Contact Process Instruments about returning the laser. If the laser appears to be working properly, please see the section on Serial control board not responding.
 
laser_power_is_reading_zero.txt · Last modified: 2008/07/09 15:18 by ejacobsen
 
Recent changes RSS feed Valid XHTML 1.0 Valid CSS Driven by DokuWiki